### v4

# TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation

Abstract
[…]
TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation
Renren Jin1thanks: Work done during internship at Xiaomi AI Lab., Tianhao Shen1, Xinwei Wu1, Dan Shi1, Haoran Sun1, Yuqi Ren1 Wuwei Huang2, Quandong Wang2, Wei Liu2, Jian Luan2, Bin Wang2, Deyi Xiong1thanks:  Corresponding author. 1School of Computer Science and Technology, Tianjin University, Tianjin, China 2Xiaomi AI Lab, Beijing, China {rrjin, thshen, wuxw2021, shidan, hrsun, ryq20, dyxiong}@tju.edu.cn {huangwuwei, wangquandong, liuwei40, luanjian, wangbin11}@xiaomi.com
1 Introduction
Large language models (LLMs) typically undergo three primary training stages: (1) pre-training, during which LLMs are trained on extensive and diverse datasets encompassing multiple languages and modalities (Zhao et al., 2023; Yang et al., 2024; Dubey et al., 2024; DeepSeek-AI et al., 2024); (2) supervised fine-tuning, in which LLMs are fine-tuned using prompt–response pairs authored by humans or LLMs (Zhang et al., 2023b); and (3) preference fine-tuning, wherein preference data annotated by humans or LLMs are used to conduct RLHF or alternative methods such as DPO and KTO (Ouyang et al., 2022; Rafailov et al., 2023; Ethayarajh et al., 2024). The latter two stages are essential for improving LLMs’ ability to follow instructions and generate responses aligned with human preferences and values.
Figure 1: Diagram of the preference dataset construction process using TaP, where the taxonomy can be derived from the undergraduate program catalog.
However, datasets for supervised and preference fine-tuning often require human annotation, which is time-consuming and expensive. These constraints hinder the scalability of dataset expansion, limiting the amount of available training data. A common alternative is to leverage LLMs to generate synthetic data, thereby reducing reliance on human labor and improving scalability. Nevertheless, naive prompting often yields homogeneous data with limited coverage and pronounced task or domain imbalance, where certain domains or tasks are overrepresented while others remain underrepresented (Wang et al., 2023; Gunasekar et al., 2023; Xu et al., 2025). Such imbalance can cause LLMs to overfit overrepresented domains or tasks, weakening their generalizability in real-world applications. Consequently, ensuring both diversity and comprehensive coverage in synthetic data generation remains a critical challenge.
Furthermore, compared with open-source datasets for supervised fine-tuning, far fewer datasets are available for preference fine-tuning (Cui et al., 2024; Wang et al., 2024d, c). Additionally, open-source datasets for both supervised and preference fine-tuning are predominantly in English. This language bias limits the global applicability of LLMs and marginalizes non-English-speaking communities, preventing them from fully benefiting from these models. Although translating English datasets into other languages is a potential solution, it can introduce translation errors and produce “translationese”, thereby compromising dataset quality (Riley et al., 2020). Given that there are over 7,000 languages worldwide, it is imperative to develop novel approaches and release open-source datasets in diverse languages to better support the latter two training stages of LLMs.
To address these challenges, we propose the Taxonomy-Guided Preference Data Generation (TaP) framework for automated, scalable preference data generation across languages. TaP leverages a predefined taxonomy to guide LLMs in producing preference data spanning diverse domains and tasks. By generating data according to the taxonomy, TaP provides fine-grained control over dataset composition and facilitates the construction of diverse datasets with broad coverage of the taxonomy categories. The TaP framework comprises four steps: (1) designing a taxonomy aligned with the target application of LLMs (e.g., for general-purpose LLMs, organizing the taxonomy around use cases such as brainstorming, open-ended question answering, and summarization); (2) prompting LLMs to generate prompts for the categories specified in the taxonomy; (3) employing multiple LLMs to generate diverse responses for each prompt; and (4) using LLMs to evaluate and rank responses according to predefined criteria. A diagram of the dataset construction process using the TaP framework is shown in Figure 1.
Motivated by the fundamental role of knowledge in decision-making, problem-solving, and adaptation to new environments, as well as by empirical findings indicating that LLMs benefit from diverse and knowledge-rich training data (Li et al., 2023; Lozhkov et al., 2024), we propose leveraging a comprehensive taxonomy, which encompasses a broad spectrum of knowledge and skills essential to human expertise, to guide the generation of such preference data using LLMs. Through careful manual analysis, we identify undergraduate program catalogs as a valuable resource for constructing such a taxonomy, as they are structured, comprehensive repositories of essential knowledge for human learning. Considering that undergraduate education constitutes a foundational component of higher education systems worldwide and is available in most countries, [^1] https://en.wikipedia.org/wiki/Undergraduate_education employing undergraduate program catalogs enhances the adaptability of the TaP framework across various languages. When such catalogs are unavailable in a given language, we discuss alternative approaches for constructing the taxonomy in Appendix A.
Given the abundance of publicly available preference datasets in English and the relative scarcity of such datasets in other languages, we apply the TaP framework to generate preference data in Chinese. Specifically, we adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China [^2] http://www.moe.gov.cn/srcsite/A08/moe_1034/s4930/202403/W020240319305498791768.pdf as the guiding taxonomy for preference data generation. Manual inspection of the generated data reveals several desirable properties: (1) rich in knowledge, (2) high information density, and (3) broad coverage across domains and tasks. An illustrative example of a prompt and corresponding response is presented in Figure 3 in Appendix B to highlight these properties. We conduct both supervised and preference fine-tuning, including PPO and DPO, on LLMs using the TaP-generated data. Experimental results demonstrate that LLMs fine-tuned with TaP-generated data outperform those trained on existing open-source datasets. Our contributions can be summarized as follows:
  • We propose the TaP framework, which enables automated, scalable preference data generation across languages and provides fine-grained control over dataset composition.
  • We apply TaP to generate high-quality Chinese preference data for supervised and preference fine-tuning (e.g., PPO, DPO, GRPO). Additionally, the annotations produced by LLMs during data generation can be used to fine-tune open-source LLMs as judges (Wang et al., 2024b; Kim et al., 2024; Zhu et al., 2025b), reducing reliance on costly proprietary APIs and the risk of data leakage.
  • We conduct extensive experiments using TaP-generated preference data to train five open-source LLMs (3B–14B) from three model families: LLaMA-3.1, Qwen2.5, and Gemma-2. The results show that models trained on TaP-generated data outperform those trained on existing open-source datasets.
2 Related Work
Our research is closely related to studies on synthetic data generation using LLMs. Pioneering studies in this field have leveraged the in-context learning capabilities of LLMs by providing demonstrations to guide the generation of synthetic samples (Honovich et al., 2023; Wang et al., 2023), typically requiring a small set of seed samples. Following this line of research, subsequent work has typically required a small number of seed samples to guide LLMs in producing synthetic samples (Xu et al., 2024; Mukherjee et al., 2023; Li et al., 2024; Yue et al., 2024; Zhu et al., 2025a; Ren et al., 2025).
In contrast, TaP uses a taxonomy to guide LLMs in autonomously generating synthetic data, eliminating the need for seed samples authored by humans or LLMs. While TaP shares the goal of constructing preference datasets with UltraFeedback (Cui et al., 2024) and the HelpSteer series (Wang et al., 2024d, c), it differs in two key aspects. First, TaP can produce prompts from scratch without relying on existing instruction datasets, enabling fine-grained control over dataset composition and improving its applicability across languages. By contrast, Cui et al. (2024) and Wang et al. (2024d, c) construct preference datasets by collecting prompts from existing instruction datasets. Second, TaP constructs Chinese preference datasets to address the scarcity of such data, whereas Cui et al. (2024) and Wang et al. (2024d, c) mainly focus on English.
3 Preference Dataset Curation
The TaP framework for automated and scalable preference data generation comprises four key steps: (1) Taxonomy Specification, which constructs a structured taxonomy to comprehensively cover common use cases of LLMs; (2) Prompt Generation, which guides LLMs to produce prompts corresponding to the categories defined in the taxonomy; (3) Response Generation and Selection, in which multiple LLMs generate responses to each prompt and a subset is selected to balance annotation cost and response diversity; and (4) Fine-grained Preference Annotation, in which LLMs evaluate and score the selected responses across multiple dimensions. Figure 1 illustrates the entire process of constructing the preference dataset using the TaP framework.
3.1 Taxonomy Specification
We adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China as the taxonomy for guiding LLMs to generate Chinese preference data. This catalog lists all undergraduate subjects offered by Chinese universities, spanning a wide range of disciplines across the humanities, social sciences, and STEM fields. It comprises 816 distinct subjects grouped into 93 discipline categories. These categories are further organized into 12 primary degree-conferring disciplines, providing a structured organization of undergraduate knowledge.
To construct preference datasets covering diverse domains and tasks, we incorporate nearly all subjects specified in the catalog into our taxonomy, except those in the Foreign Languages and Literatures discipline, which includes many low-resource languages. Instead, we selectively include 12 subjects: the top nine languages, along with Linguistics, Translation, and Business English. Consequently, our taxonomy contains 724 subjects in total: 12 from Foreign Languages and Literatures and 712 from other disciplines.
3.2 Prompt Generation
We adopt a multi-stage prompting strategy, rather than instructing LLMs to generate prompts directly, to improve both the diversity and coverage of the resulting outputs. In the first stage, LLMs are prompted to generate diverse question types and brief descriptions for each subject within a predefined taxonomy through a three-round interaction. Each round produces a distinct set of question types and descriptions, thereby increasing both the quantity and diversity of the generated outputs. The prompt used in this stage is provided in Figure 4 of Appendix C.1. In the second stage, LLMs refine the question-type descriptions from the first stage to improve their clarity and real-world relevance. This refinement aims to improve the quality of the model-generated prompts. The prompt used in this stage is shown in Figure 5 of Appendix C.1. In the third stage, we employ a two-round interaction to address the issue that LLMs occasionally produce incomplete prompts that lack essential contextual information. [^3] For instance, LLMs may generate prompts that request a summary of an article without providing the article itself. In the first round, LLMs generate a candidate prompt based on the question types and refined descriptions. In the second round, they assess whether the prompt contains all necessary information. If a prompt is identified as incomplete, the LLMs regenerate it to ensure completeness. The prompt used for this interaction is illustrated in Figure 6 of Appendix C.1. Additionally, we find that LLMs sometimes generate prompts beyond their capabilities, such as those requiring real-world actions. To mitigate this issue, we employ an LLM as a checker to evaluate whether a given prompt is feasible within its operational constraints. If a prompt is deemed infeasible, it is revised, with up to three regeneration attempts. The evaluation prompt for this step is presented in Figure 7 of Appendix C.1. To verify the reliability of LLMs in identifying such infeasible prompts, we manually reviewed a subset of evaluations generated by GPT-4. The results demonstrate that GPT-4 reliably identifies prompts that exceed its capabilities. Further details are provided in Appendix C.2.
Dataset #Samples #Prompts Prompt Length Response Length Fine-grained Annotation? Annotation Type Annotator Languages
Huozi-RLHF (Huozi-Team, 2024) 16,918 3,725 69.19 $\pm$ 107.38 216.02 $\pm$ 165.44 Ranking Human Chinese
Chinese-DPO-Pairs 10,735 10,717 220.23 $\pm$ 318.37 401.90 $\pm$ 347.63 Ranking AI Chinese
TaP (GPT-4) 261,874 27,715 379.29 $\pm$ 159.11 982.89 $\pm$ 919.97 Scalar AI Chinese
Table 1: Comparison of open-source Chinese preference datasets with the dataset constructed using TaP. “TaP (GPT-4)” indicates that the prompts in the dataset were generated by GPT-4. The column labeled “#Prompts” denotes the number of unique prompts in each dataset. “Prompt Length” and “Response Length” are reported as $XX\pm YY$, where $XX$ is the mean number of tokens and $YY$ is the standard deviation. Tokenization uses the cl100k_base encoding provided by the tiktoken library. [^5] https://github.com/openai/tiktoken
3.3 Response Generation and Selection
We select 65 widely used LLMs (0.5B–236B parameters) to generate responses for each prompt, increasing response diversity and coverage. The models span dense and MoE architectures and include proprietary and open-source models. The rationale for model selection and detailed model specifications are provided in Appendix C.3. Each model uses the system prompt shown in Figure 8 in Appendix C.1, producing 65 responses per prompt.
Because each prompt yields 65 responses, it produces 2,080 (65 $\times$ 64 / 2) pairwise comparisons for annotation. To reduce annotation cost while maintaining response diversity and coverage, we apply K-means clustering to partition the responses for each prompt into five clusters and select one representative response per cluster. Concretely, we choose the response closest to each cluster centroid, except for the cluster containing the GPT-4 response; in that case, we select the GPT-4 response to ensure that each prompt includes at least one high-quality response. For clustering, we compute response embeddings by mean-pooling token representations from the final layer of Qwen2-7B.
3.4 Fine-grained Preference Annotation
Figure 2: Distribution of prompts generated by GPT-4 (left) and DeepSeek-V2 (right) across the 12 primary disciplines defined in the taxonomy.
We employ multiple LLMs to perform preference annotations for each pairwise combination of selected responses to a given prompt. Rather than relying on a single model, each response pair is annotated by at least five different LLMs. Specifically, we use GPT-4 together with five high-performing open-source LLMs, including Qwen2-72B-Instruct, Command R+, Mistral-Large-Instruct-2407, Llama-3.1-70B-Instruct, and DeepSeek V2.5, to generate independent preference judgments for all pairwise comparisons. [^6] Due to the high cost of using GPT-4 for preference annotation, we use it only to annotate response pairs for prompts it generates. In contrast, response pairs for prompts generated by DeepSeek-V2 are annotated by open-source LLMs.
Human-1 Human-2 Human-3 GPT-4 DeepSeek V2.5 Command R+ Llama-3.1-70B-Instruct Mistral-Large-Instruct-2407 Qwen2-72B-Instruct Average
Human-1 - 78.67% 76.67% 78.93% 79.67% 77.63% 78.93% 78.19% 74.25% 77.87%
Human-2 78.67% - 80.00% 88.63% 90.33% 88.14% 88.96% 85.91% 82.94% 85.45%
Human-3 76.67% 80.00% - 83.61% 82.33% 84.07% 83.61% 84.90% 80.94% 82.02%
GPT-4 78.93% 88.63% 83.61% - 93.65% 95.25% 92.62% 93.94% 91.61% 89.78%
DeepSeek V2.5 79.67% 90.33% 82.33% 93.65% - 93.56% 95.99% 92.28% 89.63% 89.68%
Command R+ 77.63% 88.14% 84.07% 95.25% 93.56% - 92.54% 93.17% 91.84% 89.52%
Llama-3.1-70B-Instruct 78.93% 88.96% 83.61% 92.62% 95.99% 92.54% - 92.93% 89.60% 89.40%
Mistral-Large-Instruct-2407 78.19% 85.91% 84.90% 93.94% 92.28% 93.17% 92.93% - 91.58% 89.11%
Qwen2-72B-Instruct 74.25% 82.94% 80.94% 91.61% 89.63% 91.84% 89.60% 91.58% - 86.55%
Table 2: Pairwise preference annotation agreement across annotators. Each cell reports the agreement score between a pair of annotators, and the highest agreement value in each column is highlighted in bold.
For preference annotation, we adopt a pairwise comparison approach, wherein LLMs choose the preferred response between two alternatives, rather than assigning independent scores to each response as in pointwise grading. Our preliminary experiments indicate that pointwise grading often fails to distinguish subtle differences between similar responses, frequently assigning them identical scores. Despite its higher annotation cost, we employ pairwise comparison to obtain more precise preference annotations. In this approach, we present two responses to LLMs and ask them to score each one across four dimensions: relevance, correctness, clarity, and completeness. These dimensions facilitate a comprehensive assessment of response quality. The resulting scores can be used to train fine-grained reward models, which in turn enhance control over LLM-generated responses by assigning weights to rewards from different fine-grained reward models during RLHF (Wu et al., 2023). The prompt used for pairwise comparison is provided in Figure 9 of Appendix C.1. Additionally, to mitigate position bias in pairwise comparison, where LLMs tend to assign higher scores to the response appearing in a particular position (e.g., the first or second slot), we adopt the Balanced Position Calibration strategy (Wang et al., 2024a), which involves conducting an additional comparison with the positions of the two responses swapped.
To derive preference labels from pairwise comparison scores, we compute an overall score for each response and label the higher-scoring response as preferred. Specifically, we average the score for each evaluation dimension across both positions (original vs. swapped) to obtain calibrated scores. We then assign equal weights to the calibrated dimension scores and average them to produce a final score for each response. Although domain-specific applications may prioritize certain dimensions (e.g., correctness in medical applications), we use uniform weights across all four dimensions to support general-purpose LLM training.
4 Dataset Statistics
We employ GPT-4 and DeepSeek-V2 to generate prompts for the 724 subjects in our taxonomy, derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. GPT-4 generates 27,715 prompts, whereas DeepSeek-V2 produces 50,929 prompts. The distributions of these prompts across the 12 primary disciplines are presented in Figure 2. As shown in Figure 2, despite using different LLMs for prompt generation, the resulting distributions are highly consistent, indicating that TaP enables fine-grained control over dataset composition. In addition, because some prompts must be regenerated after failing the feasibility check, a single question type may correspond to multiple prompts. Consequently, GPT-4 yields 19,840 distinct question types, and DeepSeek-V2 yields 24,839 distinct question types.
For constructing the preference dataset, we rely solely on preference annotations generated by GPT-4, given its pioneering performance. [^7] While incorporating annotations from multiple LLMs might improve dataset quality, we leave the integration of annotations from other open-source LLMs for future work. Table 5 presents a comparative summary of the preference dataset constructed by TaP and existing open-source preference datasets. Compared with these datasets, the TaP-constructed dataset contains more samples and prompts, includes longer prompts and responses, and provides fine-grained annotations.
5 Annotation Agreement
To assess the reliability of preference annotations produced by LLMs, we examine three aspects of annotation agreement: (i) agreement among human annotators, (ii) agreement between human annotators and LLMs, and (iii) agreement across different LLMs. To this end, we randomly sample 300 prompts, their paired responses generated by two distinct LLMs, and the corresponding model-generated preference annotations. We then recruit three PhD students, all native Chinese speakers with research interests in computer science, to independently label the preferred response for each response pair. Label Studio [^8] https://github.com/HumanSignal/label-studio is employed to provide an annotation interface that facilitates the process. For each prompt, annotators are shown the prompt and the two LLM-generated responses (Response A and Response B) and are instructed to indicate their preference by selecting one of three options: “Response A Win,” “Response B Win,” or “Tie.”
Dataset #Samples Languages
Firefly 1,649,399 Chinese
Alpaca-GPT-4-ZH (Peng et al., 2023) 48,818 Chinese
COIG (Zhang et al., 2023a) 178,246 Chinese
MOSS-SFT (Sun et al., 2024) 1,074,551 English, Chinese
COIG-CQIA (Bai et al., 2024) 44,694 Chinese
Infinity-Instruct 757,938 Chinese
BELLE-SFT (BELLEGroup, 2023) 3,606,402 Chinese
TaP-SFT (GPT-4) 19,840 Chinese
TaP-SFT (DeepSeek-V2) 24,839 Chinese
Table 3: Comparison of open-source datasets containing a substantial number of Chinese samples for supervised fine-tuning with datasets derived from TaP-constructed preference datasets. “TaP-SFT (GPT-4)” and “TaP-SFT (DeepSeek-V2)” indicate that the prompts were generated by GPT-4 and DeepSeek-V2, respectively. For the multilingual Infinity-Instruct dataset, Chinese samples were selected based on language tags in the data.
Table 2 reports pairwise annotation agreement among different annotators on 300 randomly selected samples. Agreement among human annotators approaches 80%, exceeding the 60%–75% range commonly reported in prior work (Bai et al., 2022; Ouyang et al., 2022; Cui et al., 2024). Furthermore, agreement between human annotators and LLM annotators is close to, and in some cases exceeds, 80%, reaching 90.33%. This suggests that preference annotations generated by LLMs reliably reflect human preferences. Notably, agreement among LLM annotators consistently surpasses 90%, substantially outperforming the level of agreement observed among human annotators.
Model Dataset AlignBench MT-Bench-zh
GPT-4o DeepSeek-V3 GPT-4o DeepSeek-V3
Reasoning Language Overall Reasoning Language Overall
Qwen2.5-3B Firefly 3.22 4.34 3.78 3.38 4.37 3.87 4.41 4.45
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.16 4.87 4.51 3.99 4.75 4.37 4.83 4.68
COIG (Zhang et al., 2023a) 2.73 4.12 3.42 2.84 4.03 3.43 3.35 3.44
MOSS-SFT (Sun et al., 2024) 3.82 4.96 4.39 3.73 4.82 4.27 4.66 4.38
COIG-CQIA (Bai et al., 2024) 3.61 4.38 3.99 3.65 4.31 3.98 4.47 4.35
Infinity-Instruct 4.02 5.11 4.57 3.91 4.92 4.41 4.89 4.68
BELLE-SFT (BELLEGroup, 2023) 3.88 5.11 4.49 3.82 4.99 4.41 4.65 4.49
TaP-SFT (GPT-4) 4.99 5.85 5.42 4.57 5.46 5.02 5.79 5.53
TaP-SFT (DeepSeek-V2) 5.25 5.65 5.45 4.89 5.39 5.14 5.67 5.60
Qwen2.5-7B Firefly 3.45 4.61 4.03 3.62 4.64 4.13 4.43 4.47
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.90 5.31 5.10 4.53 5.14 4.83 5.02 4.89
COIG (Zhang et al., 2023a) 2.82 4.41 3.62 3.11 4.38 3.75 3.48 3.66
MOSS-SFT (Sun et al., 2024) 4.71 5.26 4.98 4.55 5.17 4.86 5.06 4.90
COIG-CQIA (Bai et al., 2024) 4.72 5.13 4.92 4.51 5.02 4.76 4.73 4.51
Infinity-Instruct 4.92 5.63 5.28 4.70 5.54 5.12 4.96 4.95
BELLE-SFT (BELLEGroup, 2023) 4.32 5.44 4.88 4.16 5.35 4.76 4.93 4.83
TaP-SFT (GPT-4) 6.62 6.27 6.45 6.10 5.83 5.96 6.16 5.99
TaP-SFT (DeepSeek-V2) 6.25 6.07 6.16 5.77 5.93 5.85 6.10 5.97
Llama-3.1-8B Firefly 2.40 3.97 3.19 2.40 4.07 3.24 3.72 3.69
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.77 4.28 3.53 2.64 4.12 3.38 4.21 4.28
COIG (Zhang et al., 2023a) 1.79 2.98 2.38 1.88 3.04 2.46 2.68 2.74
MOSS-SFT (Sun et al., 2024) 2.84 4.41 3.63 2.69 4.32 3.50 4.21 4.06
COIG-CQIA (Bai et al., 2024) 2.54 3.57 3.06 2.39 3.56 2.97 3.32 3.34
Infinity-Instruct 3.15 5.07 4.11 3.01 4.95 3.98 4.71 4.60
BELLE-SFT (BELLEGroup, 2023) 3.47 5.05 4.26 3.24 4.88 4.06 4.63 4.74
TaP-SFT (GPT-4) 3.87 5.14 4.50 3.40 4.73 4.06 5.24 4.95
TaP-SFT (DeepSeek-V2) 3.61 4.75 4.18 3.47 4.61 4.04 5.11 4.89
Gemma-2-9B Firefly 2.36 3.95 3.15 2.44 4.06 3.25 3.92 3.77
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.17 4.22 3.70 3.19 4.17 3.68 4.39 4.09
COIG (Zhang et al., 2023a) 1.70 2.81 2.25 1.95 2.80 2.38 2.69 2.65
MOSS-SFT (Sun et al., 2024) 2.84 4.21 3.53 2.73 4.19 3.46 4.14 4.23
COIG-CQIA (Bai et al., 2024) 2.69 3.58 3.14 2.64 3.56 3.10 3.36 3.13
Infinity-Instruct 3.72 5.21 4.47 3.57 4.99 4.28 4.78 4.45
BELLE-SFT (BELLEGroup, 2023) 3.72 5.16 4.44 3.68 5.07 4.37 4.69 4.87
TaP-SFT (GPT-4) 4.10 5.17 4.64 3.60 4.88 4.24 5.37 5.03
TaP-SFT (DeepSeek-V2) 3.96 4.90 4.43 3.61 4.58 4.10 4.90 4.69
Qwen2.5-14B Firefly 3.94 4.94 4.44 4.10 4.89 4.49 4.77 4.79
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.64 5.73 5.68 5.34 5.73 5.54 5.44 5.36
COIG (Zhang et al., 2023a) 3.30 4.84 4.07 3.38 4.86 4.12 3.63 3.56
MOSS-SFT (Sun et al., 2024) 5.25 5.59 5.42 4.96 5.48 5.22 5.26 5.04
COIG-CQIA (Bai et al., 2024) 5.14 5.27 5.20 4.94 5.26 5.10 4.92 4.78
Infinity-Instruct 5.26 5.94 5.60 5.16 5.81 5.48 5.33 5.06
BELLE-SFT (BELLEGroup, 2023) 4.70 5.70 5.20 4.55 5.61 5.08 4.99 4.83
TaP-SFT (GPT-4) 6.76 6.51 6.64 6.27 6.27 6.27 6.34 6.18
TaP-SFT (DeepSeek-V2) 6.34 5.92 6.13 6.46 6.24 6.35 6.12 6.19
Table 4: Performance of five LLMs on AlignBench and MT-Bench-zh after supervised fine-tuning on different datasets. “GPT-4o” and “DeepSeek-V3” indicate that responses were evaluated using GPT-4o and DeepSeek-V3.
When computing pairwise preference annotation agreement, we observe a small number of exceptional cases in which LLMs fail to strictly follow the annotation instructions and therefore do not produce a valid preference score, which prevents extraction of the corresponding preference label. We exclude such samples from the agreement computation. Importantly, these cases are extremely rare: among the 300 randomly selected samples, at most seven LLM-generated annotations fail to yield a valid preference label. Consequently, excluding these samples has a negligible impact on the overall results and does not affect the validity of our experimental conclusions.
6 Experiments
To assess TaP’s effectiveness and the quality of the preference datasets it constructs, we conducted supervised fine-tuning and preference fine-tuning (PPO/DPO) on LLMs using TaP-generated data.
6.1 Supervised Fine-tuning
To conduct supervised fine-tuning using the preference data generated by TaP, we utilized the prompts from the dataset along with GPT-4-generated responses. To ensure a balanced distribution of question types, we randomly selected one prompt and its corresponding response for each question type with multiple prompts. The resulting dataset comprises 19,840 samples generated by GPT-4 and 24,839 samples produced by DeepSeek-V2. A comparative summary of this supervised fine-tuning dataset, alongside other open-source Chinese datasets, is provided in Table 3.
We conducted supervised fine-tuning on five open-source LLMs: (1) Qwen2.5-3B, (2) Qwen2.5-7B, (3) Llama-3.1-8B, (4) Gemma-2-9B, and (5) Qwen2.5-14B. These LLMs span three model families, with parameter sizes ranging from 3B to 14B.
6.2 Preference Fine-tuning
We adopt two dominant approaches for preference fine-tuning using the dataset constructed by TaP: PPO and DPO. For comparison with other open-source preference datasets, we utilize Huozi-RLHF (Huozi-Team, 2024) and Chinese-DPO-Pairs [^9] https://huggingface.co/datasets/wenbopan/Chinese-dpo-pairs for preference fine-tuning. Further details on the PPO and DPO setups are provided in Appendix D.1.
Model Dataset AlignBench MT-Bench-zh
GPT-4o DeepSeek-V3 GPT-4o DeepSeek-V3
Reasoning Language Overall Reasoning Language Overall
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.02 5.23 4.63 3.88 5.04 4.46 5.03 4.94
Chinese-DPO-Pairs 4.46 5.52 4.99 4.20 5.24 4.72 5.04 4.95
TaP (GPT-4) 5.19 5.90 5.55 4.79 5.62 5.21 5.88 5.63
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.15 5.82 5.49 4.65 5.50 5.08 5.72 5.24
Chinese-DPO-Pairs 5.25 6.01 5.63 4.90 5.51 5.20 5.91 5.68
TaP (GPT-4) 5.57 6.06 5.81 4.90 5.65 5.28 6.03 5.79
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.53 5.39 4.96 4.37 5.34 4.85 5.25 4.98
Chinese-DPO-Pairs 4.88 5.53 5.21 4.64 5.40 5.02 5.23 5.23
TaP (GPT-4) 5.11 5.74 5.43 4.87 5.58 5.22 5.54 5.38
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.24 6.34 6.29 5.83 6.00 5.92 6.10 5.85
Chinese-DPO-Pairs 6.49 6.42 6.46 5.94 6.09 6.01 6.16 5.98
TaP (GPT-4) 6.59 6.66 6.62 6.03 6.18 6.11 6.36 6.11
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.52 5.19 4.36 3.29 5.06 4.17 4.85 4.58
Chinese-DPO-Pairs 3.53 5.57 4.55 3.29 5.29 4.29 4.94 4.88
TaP (GPT-4) 3.96 5.62 4.79 3.50 5.37 4.43 5.41 5.14
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.70 5.22 4.46 3.33 4.86 4.09 5.19 4.84
Chinese-DPO-Pairs 4.07 5.36 4.71 3.59 5.07 4.33 5.30 4.84
TaP (GPT-4) 4.10 5.50 4.80 3.52 4.99 4.25 5.59 5.16
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.93 5.22 4.58 3.54 5.05 4.29 4.77 4.62
Chinese-DPO-Pairs 3.86 5.32 4.59 3.61 5.21 4.41 4.91 4.84
TaP (GPT-4) 4.55 5.90 5.22 4.09 5.44 4.76 5.54 5.23
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.18 5.23 4.70 3.77 4.93 4.35 5.39 5.01
Chinese-DPO-Pairs 4.67 5.52 5.09 4.09 5.21 4.65 5.54 5.20
TaP (GPT-4) 4.72 5.61 5.16 4.05 5.14 4.60 5.78 5.33
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.75 5.90 5.82 5.64 5.81 5.73 5.41 5.48
Chinese-DPO-Pairs 5.85 5.96 5.90 5.47 5.79 5.63 5.54 5.58
TaP (GPT-4) 6.44 6.39 6.42 6.05 6.17 6.11 5.84 5.89
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.94 6.64 6.79 6.41 6.28 6.35 6.25 6.11
Chinese-DPO-Pairs 6.99 6.85 6.92 6.41 6.43 6.42 6.34 6.14
TaP (GPT-4) 7.27 7.08 7.17 6.52 6.52 6.52 6.66 6.35
Table 5: Performance comparison of LLMs trained with DPO using different datasets on AlignBench and MT-Bench-zh. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. Additionally, “GPT-4o” and “DeepSeek-V3” specify that responses were evaluated using GPT-4o and DeepSeek-V3, respectively.
Model Dataset AlignBench MT-Bench-zh
GPT-4o DeepSeek-V3 GPT-4o DeepSeek-V3
Reasoning Language Overall Reasoning Language Overall
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.13 4.88 4.51 4.44 5.36 4.90 5.25 5.20
Chinese-DPO-Pairs 5.12 5.98 5.55 4.53 5.52 5.02 5.99 5.71
TaP (GPT-4) 5.20 6.11 5.66 4.71 5.62 5.17 5.80 5.68
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.61 5.27 4.94 4.62 5.55 5.08 5.86 5.76
Chinese-DPO-Pairs 4.88 5.51 5.20 5.07 5.77 5.42 6.08 5.79
TaP (GPT-4) 5.69 6.27 5.98 4.98 5.68 5.33 6.00 5.81
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.71 5.56 5.13 4.46 5.43 4.95 5.21 5.19
Chinese-DPO-Pairs 4.41 5.33 4.87 4.31 5.52 4.91 5.23 5.14
TaP (GPT-4) 4.94 5.81 5.38 4.64 5.41 5.03 5.39 5.49
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.24 6.44 6.34 5.67 5.98 5.82 6.25 6.19
Chinese-DPO-Pairs 5.79 6.02 5.90 5.89 6.04 5.97 6.35 6.31
TaP (GPT-4) 6.29 6.62 6.46 5.73 6.15 5.94 6.36 6.38
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.72 5.27 4.50 3.41 4.97 4.19 4.93 4.89
Chinese-DPO-Pairs 2.79 4.52 3.66 3.10 5.03 4.06 4.75 4.93
TaP (GPT-4) 3.78 5.57 4.68 3.33 5.28 4.30 5.14 4.94
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.94 5.25 4.60 3.61 4.86 4.24 5.25 5.04
Chinese-DPO-Pairs 3.72 5.09 4.40 3.30 4.89 4.09 5.21 5.04
TaP (GPT-4) 4.04 5.59 4.81 3.43 4.94 4.18 5.46 5.28
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.75 4.75 4.25 3.81 5.14 4.47 4.87 4.85
Chinese-DPO-Pairs 4.05 5.48 4.77 3.73 5.25 4.49 5.03 4.99
TaP (GPT-4) 4.49 5.89 5.19 3.96 5.55 4.75 5.52 5.50
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.85 5.20 4.53 3.56 4.96 4.26 5.27 5.11
Chinese-DPO-Pairs 4.14 5.32 4.73 3.55 4.91 4.23 5.45 5.18
TaP (GPT-4) 4.42 5.58 5.00 3.91 5.23 4.57 5.63 5.26
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.61 5.90 5.75 5.69 5.93 5.81 5.78 5.78
Chinese-DPO-Pairs 6.05 6.30 6.18 5.85 6.17 6.01 5.86 5.78
TaP (GPT-4) 6.53 6.49 6.51 6.02 6.16 6.09 6.32 6.21
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.61 6.50 6.55 6.42 6.34 6.38 6.41 6.51
Chinese-DPO-Pairs 6.37 6.40 6.38 6.48 6.59 6.53 6.83 6.46
TaP (GPT-4) 7.42 7.12 7.27 6.84 6.50 6.67 6.58 6.43
Table 6: Performance comparison of LLMs trained with PPO using different datasets on AlignBench and MT-Bench-zh. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. Additionally, “GPT-4o” and “DeepSeek-V3” specify that responses were evaluated using GPT-4o and DeepSeek-V3, respectively.
6.3 Evaluation
To evaluate the performance of trained LLMs, we employ MT-Bench-zh [^10] https://github.com/HIT-SCIR/huozi/tree/main/data/mt-bench-zh as the validation set, a Chinese translation of MT-Bench (Zheng et al., 2023). For testing, we adopt AlignBench (Liu et al., 2024). Both benchmarks rely on strong LLMs (e.g., GPT-4o) to evaluate model responses. Accordingly, we report scores from both GPT-4o and DeepSeek-V3 for a comprehensive evaluation. Further details on the evaluation setup are provided in Appendix D.2.
7 Experimental Results
We present the aggregated experimental results for supervised fine-tuning, DPO, and PPO in Table 4, 5, and 6, respectively. The corresponding detailed experimental results are provided in Tables 8 through 16 in Appendix E. When trained on the same dataset, we observe that larger LLMs consistently outperform their smaller counterparts within the Qwen2.5 series. Moreover, LLMs trained on data constructed using TaP consistently demonstrate strong performance across supervised fine-tuning, DPO, and PPO. This finding underscores the effectiveness of TaP in generating high-quality training data for both supervised and preference fine-tuning of LLMs. Furthermore, the experimental results yield additional findings:
To ensure optimal model performance when scaling up data for supervised fine-tuning, it is crucial to consider factors such as data quality, diversity, and coverage.
[…]
The capabilities of pre-trained LLMs provide a critical foundation for the performance of models subsequently trained through supervised fine-tuning or preference fine-tuning.
The experimental results in Tables 4, 5, and 6 indicate that LLMs initialized from Qwen2.5-3B generally outperform those initialized from Llama-3.1-8B and Gemma-2-9B, even though the latter two have nearly three times as many parameters as Qwen2.5-3B when trained on the same dataset. We hypothesize that this gap arises from the limited volume of Chinese corpora used to pre-train Llama-3.1-8B and Gemma-2-9B, which constrains their ability to generate high-quality Chinese responses after supervised fine-tuning or preference fine-tuning. In contrast, the Qwen2.5 series of LLMs have been extensively trained on Chinese corpora, enhancing their capacity to follow instructions and better aligning with human preferences and values.
The effectiveness of applying supervised fine-tuning and preference fine-tuning using the same training dataset varies across LLMs.
The experimental results presented in Tables 4, 5, and 6 indicate that the optimal open-source datasets for fine-tuning vary across LLMs. For instance, as shown in Table 4, when GPT-4o serves as the judge, the open-source fine-tuning dataset that yields the best performance for Qwen2.5-7B on MT-Bench-zh is MOSS-SFT. However, for Llama-3.1-8B, the open-source dataset that yields the best performance under the same evaluation conditions is Infinity-Instruct. These findings align with the claim proposed by Zhang et al. (2025b), which posits that supervised fine-tuning datasets that are more closely aligned with the distribution of pre-trained LLMs tend to be more effective. Nevertheless, LLMs trained on data constructed by TaP consistently achieve strong performance and generally outperform those trained on other open-source datasets. These results demonstrates that data constructed by TaP can be effectively employed for both supervised fine-tuning and preference fine-tuning across a range of LLMs.
8 Discussion
Although LLMs trained on the synthetic dataset constructed by TaP consistently outperform those trained on other open-source datasets, the mechanisms behind this strong performance remain insufficiently understood. More broadly, while LLMs trained on different datasets can exhibit substantial performance differences, a comprehensive theory explaining why training on one dataset yields better performance than another remains an open problem, and a complete theoretical account of this phenomenon is still lacking. Nevertheless, we offer our current understanding as follows.
Recent studies have shown that LLMs benefit substantially from training data that are both diverse and knowledge-rich (Li et al., 2023; Zhang et al., 2025a). Similarly, Lozhkov et al. (2024); Allal et al. (2025) construct large-scale corpora by selectively retaining samples with high educational value. We hypothesize that the superior performance of LLMs trained on the TaP-constructed synthetic dataset, relative to those trained on open-source datasets, can be attributed to several desirable properties of the samples it contains: (1) they are knowledge-rich, (2) they exhibit high information density, and (3) they span a broad range of domains and tasks. Figure 3 in Appendix B provides an illustrative example highlighting these properties. These desirable characteristics arise primarily from the taxonomy employed in the TaP framework to guide synthetic data generation. Specifically, the taxonomy is derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China, which constitutes a comprehensive and well-structured repository of foundational knowledge across disciplines. As a result, the synthetic samples generated by TaP capture a wide spectrum of essential knowledge, thereby making them particularly effective for training LLMs.
9 Conclusion
In this paper, we propose the TaP framework to enable the automated, scalable construction of preference datasets across languages. Grounded in a structured taxonomy, TaP offers fine-grained control over dataset composition, ensuring both diversity and broad coverage. To evaluate TaP and address the scarcity of Chinese preference data, we apply TaP to construct Chinese preference datasets using a taxonomy derived from the Undergraduate Program Catalog of Regular Higher Education Institutions of China. We perform supervised fine-tuning and preference fine-tuning on LLMs ranging from 3B to 14B parameters across three model families using the TaP-generated datasets. Experimental results show that LLMs trained on TaP-generated datasets outperform those trained on open-source datasets.
Limitations
Previous studies have highlighted the risk of model collapse when training LLMs on synthetic data (Gerstgrasser et al., 2024; Kazdan et al., 2025; Shumailov et al., 2024). As the TaP framework can be employed to generate synthetic data, LLMs trained on TaP-constructed datasets may also be susceptible to model collapse; this is a limitation of TaP. Nevertheless, our work focuses on automated and scalable preference dataset construction rather than mitigating the risks associated with synthetic data. Substantial research has been devoted to preventing model collapse caused by synthetic data, such as mixing real and synthetic data (Seddik et al., 2024), selecting samples based on entropy metrics (Shi et al., 2025), and proposing new training objectives to counteract overconfidence (Shabgahi et al., 2025). Addressing these risks is beyond the scope of this study, and we leave it for future work.
Appendix A Alternative Approaches for Constructing the Taxonomy
In cases where the undergraduate program catalog is unavailable, the undergraduate program catalogs from high-resource languages can serve as representative taxonomies to guide the construction of analogous taxonomies for other languages. Furthermore, we emphasize that using the undergraduate program catalog as the taxonomy represents a specific and illustrative case of applying the TaP framework for synthetic data generation. The taxonomy employed by the TaP framework is not limited to the undergraduate program catalog. An alternative approach to constructing the taxonomy involves analyzing and categorizing the potential use cases of LLMs, a method that can be extended to other languages.
While comprehensively enumerating and categorizing all possible use cases of LLMs is inherently challenging, a practical strategy is to collect user–model interactions and analyze these conversations to systematically identify and categorize use cases (Zhao et al., 2024; Zheng et al., 2024). For low-resource languages with limited resources, a feasible approach would be to initially deploy a LLM trained on all available data for these languages, potentially supplemented with data translated from high-resource languages. Following deployment, user–model conversations can be continuously collected and analyzed to construct the taxonomy grounded in real-world use cases. Additionally, both the taxonomy and the LLMs themselves can be incrementally updated as sufficient new user-model conversations become available, thereby improving model performance over time.
Appendix B An Illustrative Example
Figure 3 presents an example prompt and its corresponding response from the preference dataset constructed using TaP. Both the prompt and the response were generated by GPT-4.
Figure 3: An example prompt and its corresponding response from the preference dataset constructed using TaP. Both the prompt and the response were generated by GPT-4. Portions of the response that were excessively long have been omitted for brevity and are indicated by blue ellipses.
Appendix C Preference Dataset Curation
C.1 Prompts
The prompts utilized in the data construction process are depicted in Figures 4 through 9.
C.2 Human Evaluation on Feasibility Check
To evaluate the reliability of LLMs in identifying instructions that exceed their capabilities, we randomly sampled 100 instructions along with the corresponding assessments generated by GPT-4 indicating whether each instruction was beyond the model’s capabilities. Sampling was performed using NumPy with a fixed random seed of 42 to ensure reproducibility.
According to GPT-4’s assessments, 71 of the 100 sampled instructions were classified as within the model’s capabilities, while the remaining 29 were identified as exceeding them. We then manually verified the correctness of these assessments. Notably, we found that 99 out of the 100 assessments were correct. This result demonstrates GPT-4’s high reliability in recognizing its own limitations when explicitly prompted with instructions that incorporate descriptions of its key limitations.
For illustrative purposes, Figure 10 presents a representative example from the sampled set, comprising both the instruction and GPT-4’s corresponding assessment. In this case, GPT-4 correctly identified the instruction as exceeding the capabilities of LLMs.
C.3 LLMs for Response Generation
To enhance the diversity and coverage of responses generated by the selected models, we aimed to include a broad spectrum of LLM families and architectures. Given the substantial computational and time costs associated with inference across such a large number of models, inference efficiency was prioritized in selecting open-source models. Specifically, aside from proprietary models such as GPT-4, inference efficiency served as the primary criterion for open-source model selection.
To facilitate efficient and scalable inference, we utilized the vLLM library, [^12] https://github.com/vllm-project/vllm which is known for its high throughput and memory efficiency in LLM inference. While vLLM supports a wide range of models, its compatibility is primarily limited to widely adopted models. Consequently, our selection was constrained to models supported by version 0.6.1.post2 of vLLM, which was the version used to generate responses for all prompts.
From the set of models compatible with this version of vLLM, we selected those that met at least one of the following criteria: (1) The model demonstrated exceptional performance across various benchmarks and attained significant popularity in academic and industrial communities (e.g., models from the LLaMA, Qwen, and Gemma families); (2) The model was trained on extensive Chinese corpora, thereby ensuring the generation of high-quality synthetic Chinese data.
Based on these criteria, we selected 63 open-source LLMs. For the remaining two models, GPT-4 (proprietary) and DeepSeek-V2 (unsupported by the vLLM version we used), we employed their official APIs to generate responses. Table 7 presents a detailed summary of the architectures and parameter sizes of all LLMs used for response generation.
Appendix D Experiments
D.1 Preference Fine-tuning
For all preference datasets used in our experiments, we randomly split each dataset into two equal subsets, using one for training reward models and the other for PPO and DPO training.
For PPO and DPO training, we initialize the policies using checkpoints from supervised fine-tuning of five open-source pre-trained LLMs. Specifically, we select the best-performing checkpoint trained on two distinct types of training data for each open-source pre-trained LLM, based on their performance on the validation dataset: (1) the highest-performing checkpoint among LLMs trained on seven open-source supervised fine-tuning datasets and (2) the highest-performing checkpoint obtained from an LLM trained on supervised fine-tuning datasets derived from our constructed preference dataset. Consequently, we conduct PPO and DPO training on a total of twelve checkpoints.
For PPO training, we first train the reward model to score the responses generated by the policy. Following common practice, we initialize the reward models from pre-trained LLMs (Gooding and Mansoor, 2023; Cui et al., 2024; Kirk et al., 2024). Specifically, the reward models are initialized from the same pre-trained LLMs used to initialize the LLMs for supervised fine-tuning, with the exception that the reward model for the policy initialized from the checkpoints derived from supervised fine-tuning of Qwen2.5-3B is initialized from Qwen2.5-14B. Additionally, we constrain the rewards produced by the reward models to a predefined range $\left[-R,R\right]$ using an auxiliary loss function to stabilize PPO training, where $R$ is a positive constant. Given a prompt $x$, a preferred response $y_{w}$ and less preferred response $y_{l}$, the reward model is parameterized by $\phi$, with $\mathcal{D}$ denoting the preference dataset used for training. The auxiliary loss function is formulated as follows:
$\mathcal{L}_{\text{aux}}\left(\phi\right)=$ $-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}$ (1)
$\quad\bigl[\max\left(0,r_{\phi}\left(x,y_{w}\right)-R\right)^{2}$
$\>\>\>\quad+\max\left(0,-r_{\phi}\left(x,y_{w}\right)-R\right)^{2}$
$\>\>\>\quad+\max\left(0,r_{\phi}\left(x,y_{l}\right)-R\right)^{2}$
$\>\>\>\quad+\max\left(0,-r_{\phi}\left(x,y_{l}\right)-R\right)^{2}\bigr]$
The auxiliary loss function is incorporated into the overall loss function for training the reward models using a hyperparameter $\lambda$ that adjusts its weight. The resulting loss function for training reward models is defined as follows:
$\mathcal{L}_{\text{RM}}\left(\phi\right)=$ $-\mathbb{E}_{\left(x,y_{w},y_{l}\right)\sim\mathcal{D}}\bigl[\lambda\mathcal{L}_{\text{aux}}\left(\phi\right)+$ (2)
$\>\>\quad\log\sigma\left(r_{\phi}\left(x,y_{w}\right)-r_{\phi}\left(x,y_{l}\right)\right)\bigr]$
The hyperparameter $\lambda$ is set to 0.05 when training reward models with our constructed preference dataset. However, due to the relatively smaller size of open-source preference datasets, we reduce $\lambda$ to 0.01 when using them for training. Additionally, the hyperparameter $R$ is fixed at 5, and all reward models are trained for one epoch, which we found sufficient for convergence.
The objective of PPO is to maximize the expected reward of responses generated by the LLM. Let the LLM policy trained using PPO be parameterized by $\theta$ and denoted as $\pi_{\theta}$, and let $\pi_{\text{ref}}$ represent the reference LLM policy. Given a KL reward coefficient $\beta$, the PPO optimization objective is defined as:
$\mathcal{J}\left(\theta\right)=$ $\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}\left(\cdot|x\right)}$ (3)
$\left[r_{\phi}\left(x,y\right)-\beta\log\left(\frac{\pi_{\theta}\left(y|x\right)}{\pi_{\text{ref}}\left(y|x\right)}\right)\right]$
For PPO training, we train the policy for two epochs when using our constructed dataset and for three epochs when using open-source datasets, given their relatively smaller size. We adjust the learning rate according to the cosine schedule with a warm-up phase of 50 steps, using a peak learning rate of 1e-6 for policy models and 5e-6 for critic models. Furthermore, we vary the KL reward coefficient within $\left\{0.01,0.05\right\}$ based on the KL divergence between the policy and reference policy during PPO training.
For DPO training, we follow a similar strategy to PPO. Specifically, we train the policy for two epochs when using our constructed dataset and for three epochs when using open-source datasets. The learning rate follows the cosine schedule with a warm-up phase of 50 step and a peak learning rate of 2e-7. Following Ivison et al. (2024), we set $\beta$ to 0.01 across all experiments.
D.2 Evaluation
For LLMs trained through supervised fine-tuning and DPO, we exclusively use scores from GPT-4o to select the best-performing checkpoint. Conversely, for models trained via supervised learning on TaP-SFT (DeepSeek-V2), we rely on DeepSeek-V3 scores due to the significant API costs associated with GPT-4o. Similarly, for models trained using PPO, we also utilize DeepSeek-V3 scores for checkpoint selection.
Appendix E Experimental Results
The detailed experimental results for supervised fine-tuning, PPO, and DPO are reported in Tables 8 through 16.
Figure 4: Prompts for generating different question types with brief descriptions of each type.
Figure 5: Prompts for refining the description of various question types.
Figure 6: Prompts for generating prompts based on specified question types and their corresponding refined descriptions.
Figure 7: Prompts for evaluating the reasonableness of a given instruction, considering the inherent limitations of LLMs.
Figure 8: Prompts used for response generation.
Figure 9: Prompts for conducting pairwise comparisons, where two responses are evaluated on a scale from 1 to 10 across four evaluation dimensions: (1) relevance, (2) correctness, (3) clarity, and (4) completeness.
Figure 10: A representative example from the sampled set in which GPT-4 correctly recognizes the instruction as exceeding the capabilities of LLMs. The figure includes both the instruction and GPT-4’s corresponding assessment.
Model Name #Parameters Model Type
gpt-4-0125-preview N/A N/A
DeepSeek-V2 236B (21B) MoE
Baichuan-13B-Chat 13B Dense
Baichuan2-7B-Chat 7B Dense
Baichuan2-13B-Chat 13B Dense
Qwen1.5-0.5B-Chat 0.5B Dense
Qwen1.5-1.8B-Chat 1.8B Dense
Qwen1.5-4B-Chat 4B Dense
Qwen1.5-7B-Chat 7B Dense
Qwen1.5-14B-Chat 14B Dense
Qwen1.5-32B-Chat 32B Dense
Qwen1.5-72B-Chat 72B Dense
Qwen1.5-110B-Chat 110B Dense
Qwen1.5-MoE-A2.7B-Chat 14B (3B) MoE
Qwen2-0.5B-Instruct 0.5B Dense
Qwen2-1.5B-Instruct 1.5B Dense
Qwen2-7B-Instruct 7B Dense
Qwen2-72B-Instruct 72B Dense
Qwen2-57B-A14B-Instruct 57B (14B) MoE
Yi-6B-Chat 6B Dense
Yi-34B-Chat 34B Dense
Yi-1.5-6B-Chat 6B Dense
Yi-1.5-9B-Chat 9B Dense
Yi-1.5-34B-Chat 34B Dense
Llama-3-8B-Instruct 8B Dense
Llama-3-70B-Instruct 70B Dense
Llama-3.1-8B-Instruct 8B Dense
Llama-3.1-70B-Instruct 70B Dense
gemma-2b-it 2B Dense
gemma-7b-it 7B Dense
gemma-1.1-2b-it 2B Dense
gemma-1.1-7b-it 7B Dense
gemma-2-2b-it 2B Dense
gemma-2-9b-it 9B Dense
gemma-2-27b-it 27B Dense
Mistral-7B-Instruct-v0.3 7B Dense
Mistral-Large-Instruct-2407 123B Dense
Mistral-Nemo-Instruct-2407 12B Dense
Mixtral-8x22B-Instruct-v0.1 141B (39B) MoE
Mixtral-8x7B-Instruct-v0.1 47B (13B) MoE
c4ai-command-r-v01 35B Dense
c4ai-command-r-plus 104B Dense
aya-23-8B 8B Dense
aya-23-35B 35B Dense
chatglm3-6b 6B Dense
glm-4-9b-chat 9B Dense
AquilaChat-7B 7B Dense
AquilaChat2-7B 7B Dense
AquilaChat2-34B 34B Dense
dbrx-instruct 132B (36B) MoE
internlm-chat-7b 7B Dense
internlm-chat-20b 20B Dense
internlm2-chat-1_8b 1.8B Dense
internlm2-chat-7b 7B Dense
internlm2-chat-20b 20B Dense
internlm2_5-1_8b-chat 1.8B Dense
internlm2_5-7b-chat 7B Dense
internlm2_5-20b-chat 20B Dense
MiniCPM-2B-dpo-bf16 2B Dense
MiniCPM-2B-sft-bf16 2B Dense
MiniCPM-1B-sft-bf16 1B Dense
Orion-14B-Chat 14B Dense
XVERSE-7B-Chat 7B Dense
XVERSE-13B-Chat 13B Dense
XVERSE-65B-Chat 65B Dense
Table 7: Model architectures and parameter counts of LLMs employed in response generation. Parameter counts enclosed in parentheses represent the number of activated parameters for MoE LLMs.
Model Dataset GPT-4o DeepSeek-V3
First turn Second turn Average First turn Second turn Average
Qwen2.5-3B Firefly 5.73 3.10 4.41 5.88 3.03 4.45
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.29 3.36 4.83 6.39 2.98 4.68
COIG (Zhang et al., 2023a) 4.18 2.53 3.35 4.56 2.31 3.44
MOSS-SFT (Sun et al., 2024) 6.31 3.00 4.66 6.01 2.74 4.38
COIG-CQIA (Bai et al., 2024) 5.99 2.95 4.47 6.01 2.69 4.35
Infinity-Instruct 6.61 3.18 4.89 6.46 2.90 4.68
BELLE-SFT (BELLEGroup, 2023) 6.11 3.18 4.65 6.04 2.94 4.49
TaP-SFT (GPT-4) 7.63 3.96 5.79 7.36 3.69 5.53
TaP-SFT (DeepSeek-V2) 7.56 3.78 5.67 7.59 3.61 5.60
Qwen2.5-7B Firefly 5.75 3.10 4.43 6.09 2.85 4.47
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.64 3.40 5.02 6.54 3.24 4.89
COIG (Zhang et al., 2023a) 4.40 2.55 3.48 4.86 2.46 3.66
MOSS-SFT (Sun et al., 2024) 6.94 3.18 5.06 6.95 2.85 4.90
COIG-CQIA (Bai et al., 2024) 6.50 2.96 4.73 6.49 2.54 4.51
Infinity-Instruct 6.61 3.31 4.96 6.96 2.94 4.95
BELLE-SFT (BELLEGroup, 2023) 6.36 3.50 4.93 6.44 3.23 4.83
TaP-SFT (GPT-4) 8.01 4.30 6.16 8.16 3.81 5.99
TaP-SFT (DeepSeek-V2) 8.39 3.81 6.10 8.49 3.45 5.97
Llama-3.1-8B Firefly 4.79 2.65 3.72 4.94 2.44 3.69
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.46 2.95 4.21 5.66 2.90 4.28
COIG (Zhang et al., 2023a) 3.39 1.98 2.68 3.63 1.86 2.74
MOSS-SFT (Sun et al., 2024) 5.28 3.15 4.21 5.35 2.76 4.06
COIG-CQIA (Bai et al., 2024) 4.36 2.26 3.32 4.66 2.03 3.34
Infinity-Instruct 6.06 3.35 4.71 6.05 3.15 4.60
BELLE-SFT (BELLEGroup, 2023) 5.98 3.27 4.63 6.30 3.18 4.74
TaP-SFT (GPT-4) 6.80 3.69 5.24 6.83 3.08 4.95
TaP-SFT (DeepSeek-V2) 6.53 3.70 5.11 6.51 3.26 4.89
Gemma-2-9B Firefly 5.00 2.84 3.92 4.98 2.56 3.77
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.56 3.21 4.39 5.51 2.66 4.09
COIG (Zhang et al., 2023a) 3.35 2.03 2.69 3.49 1.81 2.65
MOSS-SFT (Sun et al., 2024) 5.31 2.98 4.14 5.44 3.03 4.23
COIG-CQIA (Bai et al., 2024) 4.26 2.46 3.36 4.19 2.08 3.13
Infinity-Instruct 6.34 3.23 4.78 6.08 2.83 4.45
BELLE-SFT (BELLEGroup, 2023) 6.13 3.25 4.69 6.46 3.28 4.87
TaP-SFT (GPT-4) 7.03 3.71 5.37 6.89 3.16 5.03
TaP-SFT (DeepSeek-V2) 6.50 3.30 4.90 6.46 2.93 4.69
Qwen2.5-14B Firefly 6.11 3.43 4.77 6.36 3.23 4.79
Alpaca-GPT-4-ZH (Peng et al., 2023) 7.14 3.75 5.44 7.01 3.71 5.36
COIG (Zhang et al., 2023a) 4.64 2.63 3.63 4.69 2.43 3.56
MOSS-SFT (Sun et al., 2024) 6.99 3.53 5.26 7.00 3.08 5.04
COIG-CQIA (Bai et al., 2024) 6.75 3.09 4.92 6.74 2.83 4.78
Infinity-Instruct 7.00 3.66 5.33 6.93 3.19 5.06
BELLE-SFT (BELLEGroup, 2023) 6.60 3.39 4.99 6.65 3.00 4.83
TaP-SFT (GPT-4) 8.29 4.40 6.34 8.31 4.04 6.18
TaP-SFT (DeepSeek-V2) 8.20 4.04 6.12 8.29 4.09 6.19
Table 8: Performance comparison of five LLMs on MT-Bench-zh after supervised fine-tuning with various datasets. “GPT-4o” and “DeepSeek-V3” indicate that the responses were evaluated using GPT-4o and DeepSeek-V3, respectively.
Model Dataset GPT-4o DeepSeek-V3
First turn Second turn Average First turn Second turn Average
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.71 3.35 5.03 6.79 3.09 4.94
Chinese-DPO-Pairs 6.53 3.56 5.04 6.66 3.24 4.95
TaP (GPT-4) 7.69 4.08 5.88 7.76 3.49 5.63
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.46 3.98 5.72 7.18 3.31 5.24
Chinese-DPO-Pairs 7.63 4.19 5.91 7.68 3.69 5.68
TaP (GPT-4) 7.94 4.13 6.03 7.98 3.60 5.79
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.09 3.41 5.25 6.96 3.00 4.98
Chinese-DPO-Pairs 7.09 3.36 5.23 7.26 3.20 5.23
TaP (GPT-4) 7.54 3.55 5.54 7.59 3.18 5.38
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.10 4.10 6.10 7.99 3.71 5.85
Chinese-DPO-Pairs 8.04 4.29 6.16 8.24 3.73 5.98
TaP (GPT-4) 8.15 4.58 6.36 8.23 4.00 6.11
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.34 3.36 4.85 6.38 2.78 4.58
Chinese-DPO-Pairs 6.51 3.38 4.94 6.50 3.25 4.88
TaP (GPT-4) 7.11 3.70 5.41 7.11 3.16 5.14
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.81 3.56 5.19 6.59 3.09 4.84
Chinese-DPO-Pairs 6.85 3.75 5.30 6.69 2.99 4.84
TaP (GPT-4) 6.99 4.20 5.59 7.00 3.31 5.16
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.21 3.33 4.77 6.14 3.10 4.62
Chinese-DPO-Pairs 6.43 3.40 4.91 6.65 3.04 4.84
TaP (GPT-4) 7.29 3.79 5.54 7.29 3.18 5.23
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.95 3.84 5.39 6.81 3.21 5.01
Chinese-DPO-Pairs 7.00 4.09 5.54 7.00 3.40 5.20
TaP (GPT-4) 7.44 4.13 5.78 7.49 3.16 5.33
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.16 3.66 5.41 7.50 3.46 5.48
Chinese-DPO-Pairs 7.36 3.71 5.54 7.64 3.53 5.58
TaP (GPT-4) 7.64 4.05 5.84 7.81 3.96 5.89
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.21 4.29 6.25 8.30 3.93 6.11
Chinese-DPO-Pairs 8.31 4.38 6.34 8.33 3.95 6.14
TaP (GPT-4) 8.50 4.83 6.66 8.54 4.16 6.35
Table 9: Performance comparison of LLMs trained with DPO using different datasets on MT-Bench-zh. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. Additionally, “GPT-4o” and “DeepSeek-V3” specify that responses were evaluated using GPT-4o and DeepSeek-V3, respectively.
Model Dataset GPT-4o DeepSeek-V3
First turn Second turn Average First turn Second turn Average
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.86 3.64 5.25 6.91 3.49 5.20
Chinese-DPO-Pairs 7.86 4.13 5.99 7.78 3.65 5.71
TaP (GPT-4) 7.84 3.76 5.80 7.70 3.66 5.68
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.51 4.21 5.86 7.65 3.86 5.76
Chinese-DPO-Pairs 7.55 4.60 6.08 7.79 3.80 5.79
TaP (GPT-4) 7.80 4.20 6.00 7.99 3.64 5.81
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.05 3.36 5.21 7.21 3.18 5.19
Chinese-DPO-Pairs 7.10 3.36 5.23 7.03 3.25 5.14
TaP (GPT-4) 7.31 3.46 5.39 7.56 3.41 5.49
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.15 4.35 6.25 8.35 4.03 6.19
Chinese-DPO-Pairs 8.36 4.34 6.35 8.45 4.18 6.31
TaP (GPT-4) 8.54 4.19 6.36 8.55 4.21 6.38
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.34 3.53 4.93 6.46 3.33 4.89
Chinese-DPO-Pairs 6.25 3.24 4.75 6.64 3.21 4.93
TaP (GPT-4) 6.88 3.41 5.14 6.79 3.10 4.94
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.61 3.89 5.25 6.85 3.24 5.04
Chinese-DPO-Pairs 6.79 3.63 5.21 7.01 3.08 5.04
TaP (GPT-4) 6.94 3.99 5.46 6.99 3.58 5.28
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.41 3.33 4.87 6.49 3.21 4.85
Chinese-DPO-Pairs 6.58 3.48 5.03 6.64 3.34 4.99
TaP (GPT-4) 7.49 3.55 5.52 7.53 3.48 5.50
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.85 3.69 5.27 7.05 3.16 5.11
Chinese-DPO-Pairs 6.93 3.98 5.45 7.04 3.33 5.18
TaP (GPT-4) 7.25 4.01 5.63 7.29 3.23 5.26
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 7.59 3.98 5.78 7.80 3.76 5.78
Chinese-DPO-Pairs 7.80 3.93 5.86 7.81 3.74 5.78
TaP (GPT-4) 8.19 4.45 6.32 8.29 4.14 6.21
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 8.29 4.54 6.41 8.60 4.43 6.51
Chinese-DPO-Pairs 8.86 4.80 6.83 8.75 4.16 6.46
TaP (GPT-4) 8.50 4.66 6.58 8.83 4.04 6.43
Table 10: Performance comparison of LLMs trained with PPO using different datasets on MT-Bench-zh. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. Additionally, “GPT-4o” and “DeepSeek-V3” specify that responses were evaluated using GPT-4o and DeepSeek-V3, respectively.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B Firefly 3.59 2.86 3.22 4.56 4.60 4.07 4.80 3.79 4.19 4.34 3.78
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.67 3.65 4.16 4.86 4.61 4.84 4.77 5.42 4.69 4.87 4.51
COIG (Zhang et al., 2023a) 2.96 2.49 2.73 4.02 4.44 3.68 3.84 4.61 4.16 4.12 3.42
MOSS-SFT (Sun et al., 2024) 4.30 3.34 3.82 4.96 4.60 4.99 4.99 5.11 5.09 4.96 4.39
COIG-CQIA (Bai et al., 2024) 3.79 3.42 3.61 4.15 4.28 4.51 4.57 4.21 4.56 4.38 3.99
Infinity-Instruct 4.56 3.48 4.02 5.10 4.52 4.74 5.44 5.26 5.63 5.11 4.57
BELLE-SFT (BELLEGroup, 2023) 4.60 3.16 3.88 5.00 4.76 4.97 5.28 5.34 5.28 5.11 4.49
TGPDG-SFT (GPT-4) 5.92 4.05 4.99 5.59 5.50 5.26 6.28 6.39 6.07 5.85 5.42
TGPDG-SFT (DeepSeek-V2) 6.03 4.48 5.25 5.48 5.43 5.25 5.92 6.13 5.70 5.65 5.45
Qwen2.5-7B Firefly 3.66 3.24 3.45 4.72 4.31 4.97 5.07 3.84 4.78 4.61 4.03
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.49 4.30 4.90 5.50 5.21 5.44 5.24 5.13 5.34 5.31 5.10
COIG (Zhang et al., 2023a) 3.04 2.61 2.82 4.65 4.68 4.10 4.01 4.55 4.45 4.41 3.62
MOSS-SFT (Sun et al., 2024) 5.10 4.32 4.71 5.46 4.78 4.87 5.35 5.68 5.41 5.26 4.98
COIG-CQIA (Bai et al., 2024) 5.26 4.17 4.72 5.10 5.35 5.07 5.15 5.13 4.96 5.13 4.92
Infinity-Instruct 5.44 4.40 4.92 5.37 5.16 5.69 5.89 5.74 5.96 5.63 5.28
BELLE-SFT (BELLEGroup, 2023) 5.07 3.57 4.32 5.37 4.91 5.56 5.49 5.68 5.61 5.44 4.88
TGPDG-SFT (GPT-4) 7.42 5.83 6.62 6.37 5.78 5.93 6.41 6.82 6.30 6.27 6.45
TGPDG-SFT (DeepSeek-V2) 7.08 5.41 6.25 5.57 5.95 5.76 6.28 6.76 6.10 6.07 6.16
Llama-3.1-8B Firefly 2.41 2.38 2.40 3.87 3.86 3.97 4.53 3.53 4.09 3.97 3.19
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.50 3.04 2.77 4.12 3.84 4.09 4.67 4.37 4.59 4.28 3.53
COIG (Zhang et al., 2023a) 1.73 1.85 1.79 2.96 2.53 2.90 2.79 3.32 3.37 2.98 2.38
MOSS-SFT (Sun et al., 2024) 2.63 3.05 2.84 4.23 3.66 4.34 4.62 4.97 4.66 4.41 3.63
COIG-CQIA (Bai et al., 2024) 2.30 2.78 2.54 3.45 3.28 3.41 3.49 3.95 3.86 3.57 3.06
Infinity-Instruct 3.23 3.07 3.15 4.76 4.21 4.79 5.68 5.42 5.57 5.07 4.11
BELLE-SFT (BELLEGroup, 2023) 3.57 3.38 3.47 4.92 4.24 4.81 5.31 5.55 5.48 5.05 4.26
TGPDG-SFT (GPT-4) 3.83 3.90 3.87 4.90 3.83 4.85 5.52 6.05 5.71 5.14 4.50
TGPDG-SFT (DeepSeek-V2) 3.48 3.73 3.61 3.98 3.64 4.60 5.41 5.58 5.28 4.75 4.18
Gemma-2-9B Firefly 2.21 2.50 2.36 3.69 3.69 4.19 4.61 3.32 4.21 3.95 3.15
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.18 3.16 3.17 4.32 3.36 4.18 4.40 4.47 4.59 4.22 3.70
COIG (Zhang et al., 2023a) 1.47 1.92 1.70 2.94 2.38 2.40 2.56 3.26 3.30 2.81 2.25
MOSS-SFT (Sun et al., 2024) 2.72 2.96 2.84 4.00 3.38 3.89 4.55 5.16 4.30 4.21 3.53
COIG-CQIA (Bai et al., 2024) 2.55 2.83 2.69 3.25 3.03 3.76 3.53 3.87 4.05 3.58 3.14
Infinity-Instruct 3.92 3.52 3.72 5.06 4.19 5.12 5.47 5.61 5.83 5.21 4.47
BELLE-SFT (BELLEGroup, 2023) 3.97 3.47 3.72 5.11 4.33 5.26 5.36 5.53 5.34 5.16 4.44
TGPDG-SFT (GPT-4) 4.50 3.71 4.10 5.07 3.84 5.00 5.41 5.79 5.91 5.17 4.64
TGPDG-SFT (DeepSeek-V2) 4.26 3.66 3.96 4.52 4.00 4.87 5.37 5.37 5.25 4.90 4.43
Qwen2.5-14B Firefly 4.21 3.66 3.94 4.87 5.33 5.49 4.97 4.18 4.82 4.94 4.44
Alpaca-GPT-4-ZH (Peng et al., 2023) 6.14 5.13 5.64 5.94 5.53 6.12 5.37 5.84 5.59 5.73 5.68
COIG (Zhang et al., 2023a) 3.21 3.39 3.30 4.92 5.04 4.94 4.43 5.18 4.55 4.84 4.07
MOSS-SFT (Sun et al., 2024) 5.30 5.20 5.25 5.57 5.05 5.75 5.75 5.74 5.66 5.59 5.42
COIG-CQIA (Bai et al., 2024) 5.55 4.73 5.14 5.44 5.36 5.52 5.25 4.97 5.04 5.27 5.20
Infinity-Instruct 5.96 4.55 5.26 5.86 5.37 6.03 6.19 6.03 6.16 5.94 5.60
BELLE-SFT (BELLEGroup, 2023) 4.93 4.47 4.70 5.51 5.33 6.12 5.59 5.95 5.72 5.70 5.20
TGPDG-SFT (GPT-4) 7.50 6.02 6.76 6.87 5.76 6.54 6.65 6.71 6.53 6.51 6.64
TGPDG-SFT (DeepSeek-V2) 7.21 5.47 6.34 5.99 5.57 6.07 5.87 6.16 5.85 5.92 6.13
Table 11: Performance comparison of five LLMs on Alignbench after supervised fine-tuning with various datasets. The evaluation is conducted using GPT-4o, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B Firefly 3.84 2.92 3.38 4.31 4.50 4.21 4.61 4.08 4.49 4.37 3.87
Alpaca-GPT-4-ZH (Peng et al., 2023) 4.53 3.45 3.99 4.54 4.21 4.70 4.77 5.24 5.03 4.75 4.37
COIG (Zhang et al., 2023a) 2.87 2.80 2.84 3.82 3.84 3.76 3.65 4.79 4.30 4.03 3.43
MOSS-SFT (Sun et al., 2024) 4.24 3.22 3.73 4.78 4.24 4.66 5.00 5.13 5.09 4.82 4.27
COIG-CQIA (Bai et al., 2024) 3.81 3.49 3.65 3.99 4.12 4.43 4.41 4.42 4.51 4.31 3.98
Infinity-Instruct 4.45 3.38 3.91 5.01 3.93 4.50 5.45 5.16 5.45 4.92 4.41
BELLE-SFT (BELLEGroup, 2023) 4.29 3.34 3.82 4.73 4.28 4.76 5.37 5.47 5.35 4.99 4.41
TaP-SFT (GPT-4) 5.41 3.74 4.57 5.24 4.93 5.19 5.67 5.89 5.82 5.46 5.02
TaP-SFT (DeepSeek-V2) 5.58 4.20 4.89 5.15 4.86 4.90 5.75 5.84 5.82 5.39 5.14
Qwen2.5-7B Firefly 3.91 3.34 3.62 4.67 4.41 5.00 4.92 3.95 4.89 4.64 4.13
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.07 3.99 4.53 5.35 4.60 5.31 5.11 5.11 5.35 5.14 4.83
COIG (Zhang et al., 2023a) 3.35 2.87 3.11 4.57 4.24 4.13 3.87 4.79 4.70 4.38 3.75
MOSS-SFT (Sun et al., 2024) 5.09 4.01 4.55 5.30 4.47 4.85 5.16 5.76 5.50 5.17 4.86
COIG-CQIA (Bai et al., 2024) 5.12 3.90 4.51 4.98 5.09 4.81 4.96 5.26 5.01 5.02 4.76
Infinity-Instruct 4.99 4.41 4.70 5.22 5.00 5.71 5.67 5.79 5.87 5.54 5.12
BELLE-SFT (BELLEGroup, 2023) 4.73 3.59 4.16 5.10 4.74 5.59 5.37 5.68 5.64 5.35 4.76
TaP-SFT (GPT-4) 6.84 5.36 6.10 6.06 5.14 5.58 6.04 6.05 6.09 5.83 5.96
TaP-SFT (DeepSeek-V2) 6.63 4.91 5.77 5.97 5.53 5.76 6.07 6.11 6.13 5.93 5.85
Llama-3.1-8B Firefly 2.30 2.50 2.40 3.90 3.76 4.09 4.71 3.79 4.21 4.07 3.24
Alpaca-GPT-4-ZH (Peng et al., 2023) 2.35 2.93 2.64 3.95 3.50 3.88 4.49 4.34 4.56 4.12 3.38
COIG (Zhang et al., 2023a) 1.68 2.08 1.88 3.01 2.38 2.78 2.84 3.71 3.52 3.04 2.46
MOSS-SFT (Sun et al., 2024) 2.51 2.87 2.69 4.10 3.09 4.04 4.57 5.37 4.74 4.32 3.50
COIG-CQIA (Bai et al., 2024) 2.24 2.53 2.39 3.40 2.93 3.43 3.45 4.18 3.97 3.56 2.97
Infinity-Instruct 2.88 3.14 3.01 4.71 4.24 4.38 5.48 5.26 5.60 4.95 3.98
BELLE-SFT (BELLEGroup, 2023) 3.06 3.41 3.24 4.78 3.97 4.62 5.04 5.55 5.34 4.88 4.06
TaP-SFT (GPT-4) 3.31 3.48 3.40 4.73 3.14 4.46 5.32 5.34 5.37 4.73 4.06
TaP-SFT (DeepSeek-V2) 3.44 3.50 3.47 4.40 3.12 4.49 5.09 5.21 5.34 4.61 4.04
Gemma-2-9B Firefly 2.29 2.59 2.44 3.81 3.64 4.24 4.77 3.53 4.40 4.06 3.25
Alpaca-GPT-4-ZH (Peng et al., 2023) 3.18 3.21 3.19 4.24 2.76 4.25 4.47 4.74 4.59 4.17 3.68
COIG (Zhang et al., 2023a) 1.64 2.26 1.95 2.90 1.95 2.46 2.56 3.42 3.53 2.80 2.38
MOSS-SFT (Sun et al., 2024) 2.62 2.85 2.73 3.94 3.02 3.70 4.53 5.50 4.46 4.19 3.46
COIG-CQIA (Bai et al., 2024) 2.49 2.79 2.64 3.26 2.74 3.75 3.45 4.00 4.14 3.56 3.10
Infinity-Instruct 3.64 3.50 3.57 4.77 3.79 4.88 5.33 5.29 5.85 4.99 4.28
BELLE-SFT (BELLEGroup, 2023) 3.80 3.57 3.68 4.96 4.12 4.99 5.25 5.53 5.55 5.07 4.37
TaP-SFT (GPT-4) 4.00 3.20 3.60 4.84 3.33 4.76 5.25 5.26 5.81 4.88 4.24
TaP-SFT (DeepSeek-V2) 3.84 3.38 3.61 4.51 3.16 4.60 5.08 5.03 5.13 4.58 4.10
Qwen2.5-14B Firefly 4.40 3.79 4.10 4.81 5.10 5.06 5.13 4.26 4.97 4.89 4.49
Alpaca-GPT-4-ZH (Peng et al., 2023) 5.89 4.79 5.34 5.77 5.41 5.88 5.36 6.05 5.90 5.73 5.54
COIG (Zhang et al., 2023a) 3.50 3.25 3.38 4.91 4.62 4.95 4.41 5.61 4.68 4.86 4.12
MOSS-SFT (Sun et al., 2024) 5.03 4.89 4.96 5.47 4.79 5.15 5.71 5.89 5.89 5.48 5.22
COIG-CQIA (Bai et al., 2024) 5.40 4.47 4.94 5.48 5.00 5.55 5.17 5.13 5.21 5.26 5.10
Infinity-Instruct 5.71 4.61 5.16 5.81 5.09 6.09 6.01 5.71 6.12 5.81 5.48
BELLE-SFT (BELLEGroup, 2023) 4.72 4.37 4.55 5.56 4.74 6.13 5.55 5.89 5.78 5.61 5.08
TaP-SFT (GPT-4) 6.79 5.75 6.27 6.46 5.78 6.49 6.27 6.29 6.37 6.27 6.27
TaP-SFT (DeepSeek-V2) 7.36 5.57 6.46 6.37 5.81 6.63 6.17 6.24 6.22 6.24 6.35
Table 12: Performance comparison of five LLMs on Alignbench after supervised fine-tuning with various datasets. The evaluation is conducted using DeepSeek-V3, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.63 3.42 4.02 4.86 4.74 4.87 5.55 5.61 5.74 5.23 4.63
Chinese-DPO-Pairs 5.09 3.84 4.46 5.08 5.26 4.97 6.01 6.00 5.78 5.52 4.99
TaP (GPT-4) 5.78 4.60 5.19 5.71 5.25 5.27 6.51 6.29 6.41 5.90 5.55
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.01 4.29 5.15 5.64 5.49 5.37 6.23 6.21 5.98 5.82 5.49
Chinese-DPO-Pairs 6.09 4.41 5.25 5.74 5.62 5.47 6.45 6.55 6.20 6.01 5.63
TaP (GPT-4) 6.39 4.74 5.57 5.77 5.40 5.46 6.68 6.53 6.54 6.06 5.81
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.04 4.01 4.53 5.48 5.02 5.24 5.43 5.68 5.50 5.39 4.96
Chinese-DPO-Pairs 5.36 4.40 4.88 5.54 5.18 5.36 5.75 5.89 5.49 5.53 5.21
TaP (GPT-4) 5.64 4.59 5.11 5.63 5.28 5.46 6.23 5.79 6.05 5.74 5.43
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.44 5.04 6.24 6.31 5.78 6.34 6.41 6.79 6.39 6.34 6.29
Chinese-DPO-Pairs 7.46 5.53 6.49 6.52 5.91 6.15 6.64 6.79 6.53 6.42 6.46
TaP (GPT-4) 7.66 5.52 6.59 6.79 6.00 6.40 6.88 6.95 6.93 6.66 6.62
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.68 3.37 3.52 4.98 4.43 4.78 5.40 5.74 5.83 5.19 4.36
Chinese-DPO-Pairs 3.81 3.25 3.53 5.06 4.45 5.12 6.15 6.58 6.04 5.57 4.55
TaP (GPT-4) 4.25 3.67 3.96 5.13 4.72 4.94 6.40 6.03 6.48 5.62 4.79
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.47 3.93 3.70 4.84 3.93 4.87 5.71 6.18 5.78 5.22 4.46
Chinese-DPO-Pairs 4.19 3.95 4.07 5.05 3.93 4.99 6.00 6.05 6.14 5.36 4.71
TaP (GPT-4) 4.14 4.07 4.10 5.02 3.95 5.25 6.23 6.34 6.23 5.50 4.80
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.36 3.51 3.93 5.01 4.10 5.15 5.57 5.55 5.93 5.22 4.58
Chinese-DPO-Pairs 4.18 3.54 3.86 5.17 4.29 5.10 5.67 5.74 5.97 5.32 4.59
TaP (GPT-4) 4.80 4.29 4.55 5.52 4.69 5.72 6.28 6.50 6.67 5.90 5.22
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.78 3.59 4.18 4.94 3.83 5.09 5.71 5.95 5.84 5.23 4.70
Chinese-DPO-Pairs 5.12 4.22 4.67 5.23 4.00 5.46 6.11 6.11 6.23 5.52 5.09
TaP (GPT-4) 5.31 4.12 4.72 5.35 3.98 5.46 6.43 6.11 6.34 5.61 5.16
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.17 5.34 5.75 6.19 5.74 6.18 5.61 5.76 5.90 5.90 5.82
Chinese-DPO-Pairs 6.21 5.49 5.85 6.27 5.78 6.03 5.80 6.05 5.84 5.96 5.90
TaP (GPT-4) 7.11 5.78 6.44 6.67 6.19 6.38 6.24 6.39 6.44 6.39 6.42
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.54 6.35 6.94 6.95 6.38 6.75 6.55 6.68 6.51 6.64 6.79
Chinese-DPO-Pairs 7.93 6.05 6.99 7.14 6.31 6.85 6.93 7.05 6.84 6.85 6.92
TaP (GPT-4) 8.03 6.51 7.27 7.09 6.83 6.59 7.13 7.55 7.29 7.08 7.17
Table 13: Performance comparison of LLMs trained with DPO using different datasets on AlignBench. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. The evaluation is conducted using GPT-4o, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.45 3.30 3.88 4.75 4.21 4.72 5.48 5.45 5.64 5.04 4.46
Chinese-DPO-Pairs 4.69 3.71 4.20 4.77 4.83 4.71 5.67 5.66 5.82 5.24 4.72
TaP (GPT-4) 5.10 4.49 4.79 5.56 4.69 5.15 6.28 5.87 6.20 5.62 5.21
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.45 3.85 4.65 5.45 5.09 5.33 5.87 5.55 5.74 5.50 5.08
Chinese-DPO-Pairs 5.61 4.20 4.90 5.44 4.64 5.25 5.89 5.63 6.18 5.51 5.20
TaP (GPT-4) 5.79 4.02 4.90 5.27 4.88 5.28 6.40 5.95 6.14 5.65 5.28
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.86 3.88 4.37 5.52 4.66 5.19 5.43 5.84 5.40 5.34 4.85
Chinese-DPO-Pairs 5.15 4.13 4.64 5.47 4.57 5.27 5.75 5.84 5.51 5.40 5.02
TaP (GPT-4) 5.42 4.32 4.87 5.35 5.09 5.28 5.91 5.92 5.92 5.58 5.22
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.79 4.88 5.83 5.86 5.36 6.34 6.09 6.11 6.26 6.00 5.92
Chinese-DPO-Pairs 6.87 5.01 5.94 6.03 5.52 6.16 6.13 6.16 6.51 6.09 6.01
TaP (GPT-4) 6.92 5.14 6.03 6.28 5.26 6.24 6.43 6.32 6.59 6.18 6.11
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.29 3.29 3.29 4.71 3.97 4.65 5.52 5.63 5.88 5.06 4.17
Chinese-DPO-Pairs 3.49 3.10 3.29 4.91 4.05 4.87 5.85 6.00 6.08 5.29 4.29
TaP (GPT-4) 3.58 3.41 3.50 4.95 4.29 4.76 6.28 5.66 6.28 5.37 4.43
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.01 3.64 3.33 4.61 3.22 4.65 5.39 5.61 5.71 4.86 4.09
Chinese-DPO-Pairs 3.54 3.63 3.59 4.73 3.69 4.71 5.59 5.66 6.08 5.07 4.33
TaP (GPT-4) 3.46 3.58 3.52 4.52 3.26 4.49 5.73 5.82 6.10 4.99 4.25
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.72 3.35 3.54 4.81 3.72 4.99 5.49 5.39 5.90 5.05 4.29
Chinese-DPO-Pairs 3.77 3.46 3.61 5.09 3.83 5.09 5.60 5.76 5.87 5.21 4.41
TaP (GPT-4) 4.16 4.01 4.09 5.17 3.98 5.32 5.96 5.87 6.33 5.44 4.76
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.37 3.17 3.77 4.69 3.43 5.25 5.32 5.34 5.56 4.93 4.35
Chinese-DPO-Pairs 4.54 3.64 4.09 5.15 3.43 5.41 5.57 5.42 6.24 5.21 4.65
TaP (GPT-4) 4.49 3.61 4.05 4.85 3.41 5.18 5.88 5.55 5.99 5.14 4.60
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.12 5.17 5.64 5.98 5.64 5.90 5.51 5.87 5.97 5.81 5.73
Chinese-DPO-Pairs 6.05 4.89 5.47 6.04 5.43 5.88 5.63 5.87 5.91 5.79 5.63
TaP (GPT-4) 6.77 5.34 6.05 6.37 5.74 6.26 6.04 6.21 6.41 6.17 6.11
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.19 5.64 6.41 6.41 5.83 6.60 6.19 6.24 6.43 6.28 6.35
Chinese-DPO-Pairs 7.44 5.38 6.41 6.57 5.79 6.65 6.51 6.32 6.74 6.43 6.42
TaP (GPT-4) 7.31 5.73 6.52 6.43 5.91 6.63 6.77 6.42 6.96 6.52 6.52
Table 14: Performance comparison of LLMs trained with DPO using different datasets on AlignBench. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. The evaluation is conducted using DeepSeek-V3, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.62 3.65 4.13 4.45 4.71 4.57 5.03 5.18 5.36 4.88 4.51
Chinese-DPO-Pairs 5.82 4.42 5.12 5.83 5.22 5.07 6.43 6.89 6.41 5.98 5.55
TaP (GPT-4) 6.24 4.16 5.20 5.73 5.50 5.61 6.49 6.71 6.60 6.11 5.66
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.59 3.63 4.61 4.97 4.79 5.07 5.48 5.79 5.53 5.27 4.94
Chinese-DPO-Pairs 5.82 3.93 4.88 5.17 5.41 5.13 5.68 5.97 5.71 5.51 5.20
TaP (GPT-4) 6.49 4.89 5.69 5.93 6.12 5.53 6.55 6.89 6.59 6.27 5.98
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 5.04 4.39 4.71 5.60 5.00 5.68 5.61 5.92 5.53 5.56 5.13
Chinese-DPO-Pairs 5.03 3.79 4.41 5.15 5.18 5.28 5.72 5.32 5.34 5.33 4.87
TaP (GPT-4) 5.29 4.60 4.94 5.66 5.31 5.50 6.28 6.29 5.83 5.81 5.38
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 7.11 5.38 6.24 6.45 6.19 6.26 6.40 6.89 6.46 6.44 6.34
Chinese-DPO-Pairs 6.82 4.75 5.79 5.73 5.67 5.53 5.76 7.11 6.34 6.02 5.90
TaP (GPT-4) 7.05 5.53 6.29 6.77 5.91 6.35 6.81 7.13 6.75 6.62 6.46
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.78 3.66 3.72 5.02 4.36 4.91 5.67 5.92 5.74 5.27 4.50
Chinese-DPO-Pairs 2.76 2.83 2.79 4.26 3.48 4.46 4.83 4.82 5.29 4.52 3.66
TaP (GPT-4) 3.90 3.65 3.78 5.17 4.67 4.88 6.31 6.21 6.20 5.57 4.68
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.81 4.08 3.94 4.81 4.21 5.09 5.73 5.89 5.78 5.25 4.60
Chinese-DPO-Pairs 3.56 3.88 3.72 5.00 4.05 4.88 5.73 5.50 5.35 5.09 4.40
TaP (GPT-4) 3.92 4.15 4.04 5.05 4.24 5.12 6.33 6.53 6.29 5.59 4.81
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.23 3.26 3.75 5.09 3.91 4.44 4.88 5.03 5.16 4.75 4.25
Chinese-DPO-Pairs 4.38 3.73 4.05 4.89 4.63 4.97 6.15 6.24 6.01 5.48 4.77
TaP (GPT-4) 4.85 4.14 4.49 5.75 4.66 5.45 6.69 6.34 6.45 5.89 5.19
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 4.13 3.57 3.85 5.04 4.53 5.13 5.53 5.26 5.73 5.20 4.53
Chinese-DPO-Pairs 4.52 3.77 4.14 5.02 4.12 5.24 5.48 6.05 5.98 5.32 4.73
TaP (GPT-4) 4.83 4.00 4.42 5.35 4.14 5.15 6.14 6.47 6.24 5.58 5.00
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.03 5.18 5.61 5.50 6.05 6.35 5.73 5.82 5.93 5.90 5.75
Chinese-DPO-Pairs 6.60 5.51 6.05 6.58 5.81 6.46 6.27 6.42 6.24 6.30 6.18
TaP (GPT-4) 7.10 5.96 6.53 6.77 6.41 6.06 6.52 6.74 6.43 6.49 6.51
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.79 6.43 6.61 6.07 6.38 6.22 6.59 7.03 6.69 6.50 6.55
Chinese-DPO-Pairs 7.32 5.42 6.37 6.22 6.22 6.59 6.43 6.58 6.34 6.40 6.38
TaP (GPT-4) 8.07 6.76 7.42 7.30 6.74 6.81 7.19 7.42 7.27 7.12 7.27
Table 15: Performance comparison of LLMs trained with PPO using different datasets on AlignBench. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. The evaluation is conducted using GPT-4o, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.
Model Dataset Reasoning Language Overall
Math. Logi. Avg. Pro. Chi. Fund. Writ. Open. Role. Avg.
Qwen2.5-3B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.93 3.96 4.44 5.10 4.81 4.72 5.68 5.76 6.07 5.36 4.90
Chinese-DPO-Pairs 5.14 3.91 4.53 5.31 4.64 4.85 6.12 6.13 6.06 5.52 5.02
TaP (GPT-4) 5.61 3.82 4.71 5.29 4.71 5.21 6.16 5.97 6.40 5.62 5.17
Qwen2.5-3B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 5.33 3.90 4.62 5.28 4.95 5.37 5.81 5.89 6.00 5.55 5.08
Chinese-DPO-Pairs 5.83 4.30 5.07 5.57 5.43 5.25 6.27 5.92 6.16 5.77 5.42
TaP (GPT-4) 5.89 4.07 4.98 5.53 5.32 5.18 5.95 6.03 6.08 5.68 5.33
Qwen2.5-7B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 4.85 4.08 4.46 5.19 4.55 5.60 5.67 5.84 5.72 5.43 4.95
Chinese-DPO-Pairs 4.82 3.79 4.31 5.48 4.93 5.45 5.69 5.79 5.78 5.52 4.91
TaP (GPT-4) 5.03 4.25 4.64 5.39 4.84 5.09 5.68 5.82 5.66 5.41 5.03
Qwen2.5-7B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.67 4.67 5.67 6.04 5.52 5.69 6.03 6.26 6.32 5.98 5.82
Chinese-DPO-Pairs 6.87 4.92 5.89 5.97 5.62 5.85 6.20 6.29 6.34 6.04 5.97
TaP (GPT-4) 6.51 4.95 5.73 6.36 5.26 6.21 6.25 6.35 6.49 6.15 5.94
Llama-3.1-8B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.54 3.27 3.41 4.81 3.79 4.43 5.16 5.82 5.80 4.97 4.19
Chinese-DPO-Pairs 3.10 3.11 3.10 4.91 3.83 4.78 5.49 5.26 5.88 5.03 4.06
TaP (GPT-4) 3.43 3.23 3.33 4.99 4.10 4.40 6.09 5.97 6.12 5.28 4.30
Llama-3.1-8B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.50 3.73 3.61 4.28 3.66 4.53 5.36 5.50 5.81 4.86 4.24
Chinese-DPO-Pairs 3.13 3.47 3.30 4.47 3.48 4.69 5.55 5.55 5.59 4.89 4.09
TaP (GPT-4) 3.17 3.68 3.43 4.52 3.31 4.63 5.60 5.66 5.90 4.94 4.18
Gemma-2-9B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 3.98 3.63 3.81 4.98 4.00 4.99 5.25 5.92 5.71 5.14 4.47
Chinese-DPO-Pairs 4.06 3.39 3.73 4.84 4.14 4.87 5.88 5.68 6.09 5.25 4.49
TaP (GPT-4) 4.26 3.65 3.96 5.59 3.84 5.37 6.16 6.08 6.24 5.55 4.75
Gemma-2-9B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 3.62 3.50 3.56 4.69 3.83 4.96 5.17 5.50 5.61 4.96 4.26
Chinese-DPO-Pairs 3.82 3.28 3.55 4.69 3.28 4.93 5.40 5.42 5.73 4.91 4.23
TaP (GPT-4) 4.31 3.51 3.91 4.98 3.67 4.96 5.80 5.92 6.07 5.23 4.57
Qwen2.5-14B-SFT-Open Huozi-RLHF (Huozi-Team, 2024) 6.53 4.86 5.69 5.83 5.71 6.25 5.64 6.03 6.10 5.93 5.81
Chinese-DPO-Pairs 6.35 5.35 5.85 6.22 5.62 6.50 6.08 6.32 6.29 6.17 6.01
TaP (GPT-4) 6.73 5.32 6.02 6.13 5.79 6.00 6.28 6.32 6.43 6.16 6.09
Qwen2.5-14B-SFT-TaP Huozi-RLHF (Huozi-Team, 2024) 6.94 5.90 6.42 6.55 5.88 6.58 6.17 6.34 6.49 6.34 6.38
Chinese-DPO-Pairs 7.23 5.73 6.48 6.68 6.14 6.91 6.57 6.50 6.72 6.59 6.53
TaP (GPT-4) 7.38 6.29 6.84 6.67 6.05 6.35 6.64 6.45 6.81 6.50 6.67
Table 16: Performance comparison of LLMs trained with PPO using different datasets on AlignBench. The model names include two possible suffixes: “Open” and “TaP.” The “Open” suffix indicates that the LLMs were initialized from models trained via supervised fine-tuning on open-source datasets, whereas “TaP” denotes initialization from models trained on a dataset constructed by TaP. The evaluation is conducted using DeepSeek-V3, which scores the models’ responses. The “Fund.” column denotes Fundamental Language Ability, “Chi.” denotes Advanced Chinese Understanding, “Open.” denotes Open-ended Questions, “Writ.” denotes Writing Ability, “Role.” denotes Task-oriented Role Play, “Pro” denotes “Professional Knowledge”, “Math.” denotes Mathematics, and “Logic.” denotes Logical Reasoning.